OpenStudy (anonymous):

showing steps solve 2b+4=-18-9b

jhonyy9 (jhonyy9):

11b=-22 b=-22/11 b=-2

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

first step add 9b to both sides

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

will get 11b+4=-18-9b+9b+4 11b+4= -18 now subtract 4 from both sides and will get 11b+4-4= -18-4 11b=-22 b= -22/11 b= -2

jhonyy9 (jhonyy9):

so if is right please push the good answer button

jhonyy9 (jhonyy9):

thank you